The weights (in kg) of 11 Stem B follow a normal distribution and has a mean of 52 and a standard deviation of 4. How many students have weights greater than 56?

1
Expert's answer
2022-06-09T10:05:06-0400
"P(X>56)=1-P(Z\le\frac{56-52}{4})""=1-P(Z\le1)=0.158655"

"15.8655\%"


"N=11(0.158655)\approx 2."